# SpikeATac: A Multimodal Tactile Finger with Taxelized Dynamic Sensing for Dexterous Manipulation

### v1

@@ I Introduction
These types of sensing are highly complementary. While static sensors can tell us about contact forces, contact geometry, and surface features, dynamic tactile sensors can reveal the precise moment when contact is made or broken, capture slip events, and measure vibrations created from extrinsic contacts and surface textures [1, 2]. Both are an important part of rich, multimodal tactile sensation.
Given the promise of dynamic sensing, the field has dedicated important resources to transduction methods that can convert this promise into advanced manipulation abilities. In particular, polyvinylidene fluoride (PVDF) has emerged as a particularly promising material for dynamic sensing. It is highly sensitive to dynamic events, offers high frequency bandwidth, is low cost, and can be manufactured in thin, flexible, and customizable form factors with custom taxel arrangements [3, 4, 5, 6]. Furthermore, amplification circuits can even be fabricated directly on the PVDF film [7, 6], giving it a high ceiling for scaling to larger and integrated arrays.
@@ II-1 Dynamic Sensing
Examples of dynamic modalities for measuring surface vibration include event-based piezoresistive arrays [8], capacitive sensors instrumented dynamically [2], hydrophones [9], MEMS microphones [10, 11, 12, 13, 14], and accelerometers or piezoelectrics embedded in the finger surface [15, 16]. We believe that high vibration signal dimensionality and spatial distribution are valuable, especially to a learned model or policy which has the ability to parse complex, overlapping signals and time-series data [8, 17, 14]. However, scaling these sensors to high-dimensional arrays is nontrivial due in part to transducer form factors, wiring, and sampling challenges.
@@ II-2 PVDF-based Fingertips
PVDF is an established dynamic sensing modality in robotics, and previous work has shown its capability to sense contact events, slip events, and textures [2, 3, 16, 15, 18, 19, 20]. A number of works have combined PVDF strips with a static sensing modality in a fingertip [21, 19, 22, 23], however these works utilize only one or two PVDF strips, limiting their resolution. One study combined multi-taxel PVDF with a piezoresistive array [5], and another demonstrated how PVDF can be used for shear sensing when fabricated in dense multi-taxel modules [24]. Neither of these works, however, demonstrate multi-taxel PVDF in a fingertip.
@@ II-3 Multimodal Fingers
Beyond PVDF, one can find many multimodal finger designs in the literature using a variety of sensing modalities. While many are flat form factors [25, 26, 27, 28], we focus on multi-curved fingers that can sense diverse contacts during dexterous manipulation. Saloutos et al. built a spherical multimodal fingertip combining barometers and time-of-flight sensors [29], BioTac combines impedance sensing electrodes in both DC and AC configurations, temperature sensors, and a hydrophone for higher frequency vibration sensing [9], and TacTip and Minsound add a MEMS microphone to a camera-based tactile finger [10, 11]. Digit360 combines many modalities into a compact fingertip, including a camera, IMU, MEMS microphones, barometers, temperature sensors, and gas sensors [12].
All of these are valuable platforms for multimodal, dynamic manipulation. The primary innovation of our work is the exploration of multi-taxel PVDF in a multimodal finger. Notably, multi-taxel PVDF provides distributed taxel sensing across the finger, rather than only a few dynamic and very high frequency (>> 1 kHz) transducers.
@@ II-4 On-robot RL fine-tuning
RL has proven to be an effective approach for policy adaptation and improvement. A growing line of research [30, 31, 32] explores using RL to fine-tune pretrained policies directly on physical robots, thereby bridging the gap between offline training and real-world deployment. SERL [33], an open-source software suite, supports on-robot RL through classifier-based rewards and human-labeled sparse signals. DPPO [34], which fine-tunes diffusion-based policies on real hardware, demonstrates the potential of combining generative policy architectures with on-robot optimization. In this work, we integrate semi-sparse human reward labels with dense tactile-based rewards. This hybrid formulation leverages the efficiency of human guidance while grounding the learning process in rich contact information, enabling more reliable improvement of base policies during real-world fine-tuning.
@@ IV Characterization and Sensitivity Analysis
The taxelized nature of our PVDF sensor is a key feature of this work. To test its behavior, we also probe the finger in 9 different locations to evaluate spatial information contained in the PVDF signals. These indentations are $1.5\text{\,}\mathrm{N}$ and normal to the surface (via different finger mounts). A heatmap of the maximum absolute signals from each taxel (Fig. 4) shows that the PVDF provides strong spatial information. We note that the traces routed on the sample also contribute signal. For example, the indentation at position 9 (bottom right) activates the upper right taxel (Fig. 4) because the upper right taxel’s trace is routed near position 9 (see trace layout in Fig. 2).
@@ V Fast and Delicate Grasping
Figure 5: (a) Setup of the fast and delicate grasping experiment, in which we command a slow, medium, or fast velocity to the gripper and stop after contact is detected, comparing PVDF-based and capacitive-based detection (Tab. I). (b) After 15 trials, the more fragile object (seaweed) is noticeably crumpled when using capacitive-based stopping, but appears nearly untouched in the PVDF condition at fast velocity. In fact, in the medium and fast conditions, the capacitive method often failed to detect contact at all. (c) Raw signals from one finger show a distinct PVDF spike at seaweed contact, while the capacitive signal does not reliably rise above its noise floor (each line shows an individual taxel from one of the two fingers).
@@ V-A Experiment Setup
The methods above are only for the initial contact detection of either finger. In both methods, to ensure a stable grasp, we follow the contact detection with a slow movement that stops when the mean of all capacitive values passes a second, object-specific threshold. For the PVDF and capacitive methods, respectively, we use a threshold of (40, 80) and (5.5, 6.5) counts for (finger 1, finger 2). We use a capacitive threshold of 3 (seaweed) and 5 (sponge) counts for the stable contact adjustment.
@@ V-B Results
We note that we cannot rule out proximity signals caused by static charge contributing to the PVDF’s contact detection. However, signals caused by proximity tend to be small relative to contact signals (Fig. 3). The sharp edge and large magnitude in the raw PVDF data (Fig. 5) and consistent stopping behavior suggest the sensing of the contact event itself plays a critical role.
Finally, we use the fast grasping capability demonstrated here to grasp and pick up a wide set of delicate objects. See the supplementary video for examples..
@@ VI Learning Delicate Object In-hand Rotation
So far, we have focused on standalone characterization of SpikeATac, as well as its integration with a parallel gripper for fast yet delicate manipulation. However, a traditional limitation of complex tactile sensors is the difficult integration in fully dexterous, multifingered manipulation. This is particularly true for current data-driven sensorimotor policies, for example for in-hand object re-orientation, which often train in simulation [36, 37, 38]: if a sensor provides highly complex signals which are difficult to simulate, how can it be used for learned motor control?
@@ VI-A On-Robot RL Fine-tuning for Behavior Improvement
While $\pi_{\mathtt{IL}}$ works well on hard objects (shown in our supplementary video), it immediately fails on fragile ones, as it has been trained with rigid objects and lacks the ability to produce delicate touch. We thus fine-tune it using RL, specifically Soft Actor-Critic [39], on raw sensor data acquired from our sensor. This pipeline shows how the highly-sensitive, difficult-to-simulate sensor signals can be integrated into learning-based manipulation.
@@ VI-B Observation and Action Space
The observation contains the length 64 history buffer of sensor signals (16 PVDF signals and 7 capacitive signals per each of 4 fingers) and the 12 joint encoder readings. The action space is the hand’s 12 joint position set points for its 12 degrees of freedom (1 roll joint and 2 flexion/extension joints per finger).
@@ VI-C Results
We compare RL fine-tuning results to an IL baseline that is trained on all available demonstration data after all fine-tuning iterations, with the IL baseline achieving 5.7 radians rotation and 0.7 destruction rate, notably lower than the RL fine-tuning results. We emphasize that the data we use for the IL fine-tuning baseline is usually hard to obtain in an IL setting, given the quick failures of the base policy. Here, we filter the RL fine-tuning dataset using human labels and only use the good trajectories for imitation learning.
These experiments show that this system and pipeline as a whole enable this difficult, delicate task. We note that the resulting policy in these experiments may be overfit to these particular objects and object sizes, and that we also don’t have a definitive analysis of what each sensor contributes. Our primary takeaway, however, is that this pipeline, which is designed for the real robot and is using real sensor signals in the reward, allows effective use of SpikeATac data for new data-driven capabilities, despite the sensor being, as of now, very difficult to simulate.
Figure 8: Average rotation amount and object destruction rate with each RL fine-tuning learning iteration evaluated on a paper hexagonal prism and cylinder (figure shows combined results). The images show examples of objects after manipulation.
@@ VII Discussion and Conclusions
A differential amplifier design and more advanced shielding and mixed signal design will likely improve the PVDF’s performance [40], which we plan to investigate in future work. We note that in addition to sensitivity to static charge, our sensor is also sensitive to how the system is grounded, but we found that with proper grounding the sensor signals were robust enough to enable all the experiments presented here.
@@
Overall, we believe SpikeATac is a powerful platform for studying multimodal tactile sensing and manipulation. The sensor’s many-taxel PVDF enables fast and delicate grasping. In conjunction with on-robot RL fine-tuning, we can integrate its raw, difficult-to-simulate multimodal sensor signals into complex learning-based manipulation policies that show unprecedented capabilities, such as in-hand reorientation of fragile objects. We believe that the combination of static and dynamic tactile sensing, integrated into capable hands and used in conjunction with powerful sensorimotor learning methods, shows great promise for future advances in robot dexterity.

### v2

@@ I Introduction
These types of sensing are highly complementary. While static sensors can tell us about contact forces, contact geometry, and surface features, dynamic tactile sensors can reveal the precise moment when contact is made or broken, capture slip events, and measure vibrations created from extrinsic contacts and surface textures [johansson_coding_2009, cutkosky_dynamic_2014]. Both are an important part of rich, multimodal tactile sensation.
Given the promise of dynamic sensing, the field has dedicated important resources to transduction methods that can convert this promise into advanced manipulation abilities. In particular, polyvinylidene fluoride (PVDF) has emerged as a particularly promising material for dynamic sensing. It is highly sensitive to dynamic events, offers high frequency bandwidth, is low cost, and can be manufactured in thin, flexible, and customizable form factors with custom taxel arrangements [kappassov_tactile_2015, bae_pvdf-based_2019, yu_tri-modal_2018, hsu_locally_2011]. Furthermore, amplification circuits can even be fabricated directly on the PVDF film [dahiya_design_2009, hsu_locally_2011], giving it a high ceiling for scaling to larger and integrated arrays.
@@ II-1 Dynamic Sensing
Examples of dynamic modalities for measuring surface vibration include event-based piezoresistive arrays [taunyazov_extended_2021], capacitive sensors instrumented dynamically [cutkosky_dynamic_2014], hydrophones [fishel_bayesian_2012], MEMS microphones [pestell_artificial_2022, andrussow_adding_2025, lambeta_digitizing_2024, chang_investigation_2024, yang_touch_2023], and accelerometers or piezoelectrics embedded in the finger surface [howe_dynamic_1993, son_tactile_1994]. We believe that high vibration signal dimensionality and spatial distribution are valuable, especially to a learned model or policy which has the ability to parse complex, overlapping signals and time-series data [taunyazov_extended_2021, piacenza_sensorized_2020, yang_touch_2023]. However, scaling these sensors to high-dimensional arrays is nontrivial due in part to transducer form factors, wiring, and sampling challenges.
@@ II-2 PVDF-based Fingertips
PVDF is an established dynamic sensing modality in robotics, and previous work has shown its capability to sense contact events, slip events, and textures [cutkosky_dynamic_2014, kappassov_tactile_2015, son_tactile_1994, howe_dynamic_1993, howe_grasping_1990, shi_surface_2023, guo_vibrotact_2024]. A number of works have combined PVDF strips with a static sensing modality in a fingertip [Choi_development_2006, shi_surface_2023, Goger_tactile_2009, ke_fingertip_2019], however these works utilize only one or two PVDF strips, limiting their resolution. One study combined multi-taxel PVDF with a piezoresistive array [yu_tri-modal_2018], and another demonstrated how PVDF can be used for shear sensing when fabricated in dense multi-taxel modules [yu_flexible_2016]. Neither of these works, however, demonstrate multi-taxel PVDF in a fingertip.
@@ II-3 Multimodal Fingers
Beyond PVDF, one can find many multimodal finger designs in the literature using a variety of sensing modalities. While many are flat form factors [lu_gtac_2022, lima_dynamic_2020, martinez-hernandez_soft_2023, romano_human-inspired_2011], we focus on multi-curved fingers that can sense diverse contacts during dexterous manipulation. Saloutos et al. built a spherical multimodal fingertip combining barometers and time-of-flight sensors [saloutos_design_2023], BioTac combines impedance sensing electrodes in both DC and AC configurations, temperature sensors, and a hydrophone for higher frequency vibration sensing [fishel_bayesian_2012], and TacTip and Minsound add a MEMS microphone to a camera-based tactile finger [pestell_artificial_2022, andrussow_adding_2025]. Digit360 combines many modalities into a compact fingertip, including a camera, IMU, MEMS microphones, barometers, temperature sensors, and gas sensors [lambeta_digitizing_2024].
All of these are valuable platforms for multimodal, dynamic manipulation. The primary innovation of our work is the exploration of multi-taxel PVDF in a multimodal finger. Multi-taxel PVDF provides distributed sensing across the finger, rather than only a few dynamic, high frequency (>> 1 kHz) transducers. Lastly, we note that high bandwidth static sensors can also capture high frequency information. We favor a multimodal approach with designated dynamic sensors like PVDF so that the dynamic modality can be tuned to be sensitive to small transient events, while static sensors can preserve a larger dynamic range.
@@ II-4 On-robot RL fine-tuning
RL has proven to be an effective approach for policy adaptation and improvement. A growing line of research [smith2024grow, smith2022walk, zhou2024efficient] explores using RL to fine-tune pretrained policies directly on physical robots, thereby bridging the gap between offline training and real-world deployment. SERL [luo2024serl], an open-source software suite, supports on-robot RL through classifier-based rewards and human-labeled sparse signals. DPPO [dppo2024], which fine-tunes diffusion-based policies on real hardware, demonstrates the potential of combining generative policy architectures with on-robot optimization. In this work, we integrate semi-sparse human reward labels with dense tactile-based rewards. This hybrid formulation leverages the efficiency of human guidance while grounding the learning process in rich contact information, enabling more reliable improvement of base policies during real-world fine-tuning.
@@ IV Characterization and Sensitivity Analysis
The taxelized nature of our PVDF sensor is a key feature of this work. To test its behavior, we also probe the finger in 9 different locations to evaluate spatial information contained in the PVDF signals. These indentations are $1.5\text{\,}\mathrm{N}$ and normal to the surface (via different finger mounts). A heatmap of the maximum absolute signals from each taxel (Fig. 4) shows that the PVDF provides spatial contact information. We note that the traces routed on the sample also contribute signal. For example, the indentation at position 9 (bottom right) activates the upper right taxel because the upper right taxel’s trace is routed near position 9 (see trace layout in Fig. 2). We note that position 5 has a weaker response due to a disconnected trace in the finger used for testing.
@@ V Fast and Delicate Grasping
Figure 5: (a) Setup of the fast and delicate grasping experiment, in which we command a slow, medium, or fast velocity to the gripper and stop after contact is detected, comparing PVDF-based and capacitive-based detection (Tab. I). (b) After 15 trials, the more fragile object (seaweed) is noticeably crumpled when using capacitive-based stopping, but appears nearly untouched in the PVDF condition at fast velocity. In fact, in the medium and fast conditions, the capacitive method often failed to detect contact at all. (c) Raw signals from one finger show a distinct PVDF spike at seaweed contact, while the capacitive signal does not reliably rise above its noise floor (each line/color shows an individual taxel from one of the two fingers).
@@ V-A Experiment Setup
The methods above are only for the initial contact detection of either finger. In both methods, to ensure sufficient grasp force for lifting the object, we follow the contact detection with a slow movement that stops when the mean of all capacitive values passes a second, object-specific threshold. For the PVDF and capacitive methods, respectively, we use a threshold of (40, 80) and (5.5, 6.5) counts for (finger 1, finger 2). We use a capacitive threshold of 3 (seaweed) and 5 (sponge) counts for the stable contact adjustment.
@@ V-B Results
We note that we cannot rule out proximity signals caused by static charge contributing to the PVDF’s contact detection. However, signals caused by proximity tend to be small relative to contact signals (Fig. 3), at least for the objects investigated in these experiments. The sharp edge and large magnitude in the raw PVDF data (Fig. 5) and consistent stopping behavior suggest the sensing of the contact event itself plays a critical role.
Finally, we use the fast grasping capability demonstrated here to grasp and pick up a wide set of delicate objects. See the supplementary video for examples.
@@ VI Learning Delicate Object In-hand Rotation
So far, we have focused on standalone characterization of SpikeATac, as well as its integration with a parallel gripper for fast yet delicate manipulation. However, a traditional limitation of complex tactile sensors is the difficult integration in fully dexterous, multifingered manipulation. This is particularly true for current data-driven sensorimotor policies, for example for in-hand object re-orientation, which often train in simulation [khandate_r_2024, wang_lessons_2024, qi_general_2023]: if a sensor provides highly complex signals which are difficult to simulate, how can it be used for learned motor control?
@@ VI-A On-Robot RL Fine-tuning for Behavior Improvement
While $\pi_{\mathtt{IL}}$ works well on hard objects (shown in our supplementary video), it immediately fails on fragile ones, as it has been trained with rigid objects and lacks the ability to produce delicate touch. We thus fine-tune it using RL, specifically Soft Actor-Critic [haarnoja2018soft], on raw sensor data acquired from our sensor. This pipeline shows how the highly-sensitive, difficult-to-simulate sensor signals can be integrated into learning-based manipulation.
@@ VI-B Observation and Action Space
The observation contains the length 64 history buffer of sensor signals (16 PVDF and 7 capacitive signals per finger) and the 12 joint encoder readings. The action space is the hand’s 12 joint position set points for its 12 degrees of freedom (1 roll joint and 2 flexion/extension joints per finger).
@@ VI-C Results
We compare RL fine-tuning results to an IL baseline that is trained on all available demonstration data after all fine-tuning iterations, with the IL baseline achieving 5.7/2.9 (mean/mdn.) radians rotation and 0.7 destruction rate, notably lower than the RL fine-tuning results. We emphasize that the data we use for the IL fine-tuning baseline is usually hard to obtain in an IL setting, given the quick failures of the base policy. Here, we filter the RL fine-tuning dataset using human labels and only use the good trajectories for imitation learning.
These experiments show that this system and pipeline as a whole enable this difficult, delicate task. We note that the resulting policy in these experiments may be overfit to these particular objects and object sizes, and that we also don’t have a definitive analysis of what each sensor contributes. Our primary takeaway, however, is that this pipeline, which is designed for a real robot using real sensor signals in the reward, allows the use of SpikeATac data for new data-driven capabilities, despite the sensor being, as of now, very difficult to simulate.
Figure 8: Mean and median rotation amount and object destruction rate (# times crushed out of 10) with each RL fine-tuning learning iteration evaluated on a paper hexagonal prism and cylinder (figure shows combined results). The images show examples of objects after manipulation at each iteration. The shaded region shows the interquartile range (IQR).
@@ VII Discussion and Conclusions
A differential amplifier design and more advanced shielding will likely improve the PVDF’s performance [yeiser_umbomic_2024]. To improve connector robustness, we’d also like to eliminate the ACF by connecting PVDF electrodes directly to a flexPCB [yeiser_umbomic_2024]. We note that in addition to sensitivity to static charge, our sensor is also sensitive to how the system is grounded, but we found that with proper grounding the sensor signals were robust enough to enable all the experiments presented here.
@@
Overall, we believe SpikeATac is a powerful platform for studying multimodal tactile sensing and manipulation. The sensor’s many-taxel PVDF enables fast and delicate grasping. In conjunction with on-robot RL fine-tuning, we can integrate its raw, difficult-to-simulate multimodal sensor signals into complex learning-based policies that show unprecedented capabilities, such as in-hand reorientation of fragile objects. We believe that the combination of static and dynamic tactile sensing, integrated into capable hands and used in conjunction with powerful sensorimotor learning methods, shows great promise for future advances in robot dexterity.